OpenStudy (anonymous):

y = 1/5x - 3 y = -5x +6

OpenStudy (anonymous):

-5x+6=1/5x-3 subtract six to the side -5x=1/5x-9 Divide by 1/5 on the other side -25x=9 divide by -25 -25x/-25=9/-25 x=-9/25 or -3/5

OpenStudy (anonymous):

how would i graph it??

OpenStudy (anonymous):

got it